# GRU: Mitigating the Trade-off between Unlearning and Retention for Large Language Models

### v1

@@ Appendix A Derivations and Proofs
$\Delta$ $=\mathcal{R}(\boldsymbol{\theta}_{\mathrm{u}}^{(t+1)})\;-\;\mathcal{R}(\boldsymbol{\theta}_{\mathrm{gru}}^{(t+1)})$
$\geq\Bigl{[}\underbrace{\mathcal{R}(\boldsymbol{\theta}^{(t)})\;-\;\texttt{lr}\,\langle\boldsymbol{g}_{\mathrm{r}}^{(t)},\,\boldsymbol{g}_{\mathrm{u}}^{(t)}\rangle\;+\;\frac{\ell\,\texttt{lr}^{2}}{2}\,\|\boldsymbol{g}_{\mathrm{u}}^{(t)}\|^{2}}_{\text{Lower bound for }\mathcal{R}(\boldsymbol{\theta}_{\mathrm{u}}^{(t+1)})}\Bigr{]}\;-\;\Bigl{[}\underbrace{\mathcal{R}(\boldsymbol{\theta}^{(t)})\;-\;\texttt{lr}\,\langle\boldsymbol{g}_{\mathrm{r}}^{(t)},\,\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}\rangle\;+\;\frac{L\,\texttt{lr}^{2}}{2}\,\|\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}\|^{2}}_{\text{Upper bound for }\mathcal{R}(\boldsymbol{\theta}_{\mathrm{gru}}^{(t+1)})}\Bigr{]}.$
@@
$\Delta$ $\;\geq\;\Bigl{[}-\,\texttt{lr}\,\langle\boldsymbol{g}_{\mathrm{r}}^{(t)},\,\boldsymbol{g}_{\mathrm{u}}^{(t)}\rangle\;+\;\tfrac{\ell\,\texttt{lr}^{2}}{2}\,\|\boldsymbol{g}_{\mathrm{u}}^{(t)}\|^{2}\Bigr{]}\;-\;\Bigl{[}-\,\texttt{lr}\,\langle\boldsymbol{g}_{\mathrm{r}}^{(t)},\,\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}\rangle\;+\;\tfrac{L\,\texttt{lr}^{2}}{2}\,\|\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}\|^{2}\Bigr{]}$
$\;=\;\underbrace{-\,\texttt{lr}\,\langle\boldsymbol{g}_{\mathrm{r}}^{(t)},\,\boldsymbol{g}_{\mathrm{u}}^{(t)}\rangle\;+\;\texttt{lr}\,\langle\boldsymbol{g}_{\mathrm{r}}^{(t)},\,\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}\rangle}_{\text{(linear-difference term)}}\;+\;\underbrace{\frac{\ell\,\texttt{lr}^{2}}{2}\,\|\boldsymbol{g}_{\mathrm{u}}^{(t)}\|^{2}\;-\;\frac{L\,\texttt{lr}^{2}}{2}\,\|\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}\|^{2}}_{\text{(quadratic-difference term)}}.$
@@
    Rectification Nonnegativity. Since $\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}$ is formed from $\boldsymbol{g}_{\mathrm{u}}^{(t)}$ by removing negatively aligned components with respect to $\boldsymbol{g}_{\mathrm{r}}^{(t)}$, we have $\langle\boldsymbol{g}_{\mathrm{r}}^{(t)},\,\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}\rangle\geq\langle\boldsymbol{g}_{\mathrm{r}}^{(t)},\,\boldsymbol{g}_{\mathrm{u}}^{(t)}\rangle$, and thus $-\,\langle\boldsymbol{g}_{\mathrm{r}}^{(t)},\,\boldsymbol{g}_{\mathrm{u}}^{(t)}\rangle\;+\;\langle\boldsymbol{g}_{\mathrm{r}}^{(t)},\,\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}\rangle\geq 0$. Multiplication by lr preserves non-negativity, ensuring that the expression inside the first bracket remains non-negative.
@@
    Curvature Conditions. Due to conditions a)-b), we have $\ell\;\geq\;\frac{\|\boldsymbol{g}_{\mathrm{u}}^{(t)}-\boldsymbol{g}_{\mathrm{r}}^{(t)}\|^{2}}{\|\boldsymbol{g}_{\mathrm{u}}^{(t)}+\boldsymbol{g}_{\mathrm{r}}^{(t)}\|^{2}}L$ and $\ell\;\geq\;\frac{\|\boldsymbol{g}_{\mathrm{u}}^{(t)}-\boldsymbol{g}_{\mathrm{r}}^{(t)}\|^{2}}{\|\boldsymbol{g}_{\mathrm{u}}^{(t)}+\boldsymbol{g}_{\mathrm{r}}^{(t)}\|^{2}}L+\frac{2}{\texttt{lr}}$, which further indicate that $\tfrac{\ell\,\texttt{lr}^{2}}{2}\,\|\boldsymbol{g}_{\mathrm{u}}^{(t)}\|^{2}\;-\;\tfrac{L\,\texttt{lr}^{2}}{2}\,\|\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}\|^{2}~{}\geq~{}0$, so the second bracket is also non-negative.
Thus we complete the proof. ∎
@@ Appendix B Detailed Results
This section provides comprehensive results that echo our main experiments discussed in Section 5.1. It encompasses TOFU, WMDP, and MUSE benchmarks, further incorporating additional baseline methods like SimNPO, and other metrics, such as VerbMem for MUSE. These results are summarized in Tables 6-8.
@@
Appendix C Hyper-parameter Configurations
In addition to our main results, we further discuss about our hyper-parameter configurations. We employ grid search on validation data to select proper hyper-parameters for GRU and TRU. For GRU, the candidate values for $\gamma$ include $\{0.01,0.05,0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,0.99\}$, and that for $\tau$ are $\{0.001,0.005,0.01,0.1,1.0,10,100\}$. For TRU, we select stg from the space $\{0.50,0.60,0.65,0.70,0.75,0.80,0.85\}$ while fix $\texttt{lr}=1e^{-4}$. We further summarize their detailed configurations across different setups as follows.
@@ Appendix C Hyper-parameter Configurations
Appendix D Other Analyses
@@
D.1 Visualization of Rectification.
@@
We further examine the angles between $\boldsymbol{g}_{\mathrm{u}}^{(t)}$ and $\boldsymbol{g}_{\mathrm{r}}^{(t)}$, along with those results after being rectified via GRU. We monitor the dynamics of these angles throughout the unlearning processes for various baseline methods, as well as the changes after applying GRU for rectification. As a case study, Figure LABEL:fig:cos_theta_comparison shows these results on TOFU 5% unlearning with Llama-2-7B as the backbone. Without the use of GRU, the cosine similarity between $\boldsymbol{g}_{\mathrm{u}}^{(t)}$ and $\boldsymbol{g}_{\mathrm{r}}^{(t)}$ keeps negative throughout unlearning, suggesting potential adverse effects on the overall performance of the model. In comparison, within the unlearning dynamics facilitated by GRU, it is observed that although the angles initially continue to be negative (dotted lines), our rigorous method of gradient rectification will adjust the resulting cosine similarity to exactly 0. This adjustment ensures that the gradient direction associated with unlearning is completely orthogonal to that of retention, thereby effectively maintaining the overall model performance.
D.2 Further Strategy
We further introduce a strategy to mitigate potential side effects when the unlearning and retention gradients become nearly antiparallel. We first describe the method in detail and then present its tuning results, demonstrating its impact on unlearning stability and overall model performance.
Angle Clipping for Gradient Updates. In cases where the unlearning gradient $\boldsymbol{g}_{\mathrm{u}}^{(t)}$ and the retention gradient $\boldsymbol{g}_{\mathrm{r}}^{(t)}$ are nearly antiparallel (i.e., $\cos(\boldsymbol{g}_{\mathrm{u}}^{(t)},\boldsymbol{g}_{\mathrm{r}}^{(t)})\approx-1$), directly applying our gradient adjustment from Eq. (8) can result in an extreme directional shift. Consequently, the magnitude of the updated gradient $\tilde{\boldsymbol{g}}_{\mathrm{u}}^{(t)}$ may shrink excessively, halting meaningful progress on unlearning objectives. This issue effectively stalls model updates and compromises the desired unlearning trajectory. To address the aforementioned risk of vanishing updates, we impose a constraint by setting a minimal-allowed value $c$ on $\cos(\boldsymbol{g}_{\mathrm{u}}^{(t)},\boldsymbol{g}_{\mathrm{r}}^{(t)})$. Specifically, when $\cos(\boldsymbol{g}_{\mathrm{u}}^{(t)},\boldsymbol{g}_{\mathrm{r}}^{(t)})<c$, we clip the cosine term such that:
$\cos(\boldsymbol{g}_{\mathrm{u}}^{(t)},\boldsymbol{g}_{\mathrm{r}}^{(t)})\leftarrow\max\bigl{(}\cos(\boldsymbol{g}_{\mathrm{u}}^{(t)},\boldsymbol{g}_{\mathrm{r}}^{(t)}),\,c\bigr{)}.$
By preventing $\cos(\boldsymbol{g}_{\mathrm{u}}^{(t)},\boldsymbol{g}_{\mathrm{r}}^{(t)})$ from dropping below $c$, the updated gradient avoids undesirably large rotations. This ensures more stable and controlled progress for GRU-based methods in an unlearning setup.
Tuning Results. To investigate how angle clipping affects both unlearning and retention, we performed a series of experiments on the TOFU dataset under a 5% unlearning setup (Table 9). We varied $c$ across several negative thresholds—$-0.08$, $-0.1$, $-0.4$, $-0.5$, $-0.6$, $-0.8$—and compared these against a baseline with no clipping (N/A). We evaluated performance based on two metrics: FQ and MU. Our observations suggest that moderately negative thresholds (e.g., $-0.4$ to $-0.5$) tend to strike a better balance, offering stable unlearning progress without overly compromising retention. When $c$ is set to a small negative value (e.g., $-0.08$), the resulting gradients can become too constrained, limiting the model’s ability to forget. By contrast, highly negative $c$ (e.g., $-0.8$) may trigger abrupt directional changes, potentially eroding retention. In practice, when stronger unlearning is prioritized, one may opt for a more negative threshold, whereas settings closer to 00 can help preserve model utility. Users are encouraged to tune $c$ on a validation set, selecting the threshold that appropriately balances unlearning effectiveness with overall model performance.
Table 9: Hyper-parameter turning of $c$ on TOFU with Llama-2-7B, under the 5% unlearning setup. The notation “–” indicates that the associated result is same to those without GRU.
Method Metric $c$
-0.08 -0.1 -0.4 -0.5 -0.6 -0.8 N/A
GA w/ GRU FQ $\uparrow$ -4.7631 -5.4480 -6.9414 -8.1703 -7.7503 -7.3407 -8.6008
MU $\uparrow$ 0.4300 0.4362 0.5194 0.5420 0.5667 0.5725 0.5810
NPO w/ GRU FQ $\uparrow$ -8.1703 -6.1741 -12.4230 -8.6008 -9.0417 -9.0417
MU $\uparrow$ 0.4307 0.3621 0.5396 0.5207 0.5284 0.5220
GD w/ GRU FQ $\uparrow$ -13.4850 -12.9485 -12.4230
MU $\uparrow$ 0.5624 0.5637 0.5637
Appendix E TRU Algorithm Implementation
We present the implementation details of our TRU algorithm in Algorithm 2. TRU removes knowledge from a targeted dataset without mini-batches or iterative accumulations. Specifically, it processes each data point $\boldsymbol{s}_{\mathrm{u}}$ in the unlearning set $\mathcal{D}_{\mathrm{u}}$ just once. For each data point, TRU computes a task vector $\boldsymbol{T}_{\boldsymbol{s}_{\mathrm{u}}}$ to capture the shift in parameters needed to learn $\boldsymbol{s}_{\mathrm{u}}$ from the original model $\boldsymbol{\theta}_{\mathrm{org}}$. Then, rather than relying on a gradient over $\mathcal{D}_{\mathrm{u}}\setminus\{\boldsymbol{s}_{\mathrm{u}}\}$, the retain step refines the model by computing a separate retain task vector $\boldsymbol{T}_{\mathrm{ret}}$ through finetuning on the remaining unlearning samples, indicating how to preserve overall performance. If negating $\boldsymbol{T}_{\boldsymbol{s}_{\mathrm{u}}}$ conflicts with $\boldsymbol{T}_{\mathrm{ret}}$, we remove the problematic directional component via rectification, yielding a rectified vector $\tilde{\boldsymbol{T}}_{\boldsymbol{s}_{\mathrm{u}}}$. We accumulate these rectified vectors over all points in $\mathcal{D}_{\mathrm{u}}$, and finally apply a single large-scale update to subtract the aggregate unlearning direction from the original parameters, scaled by a hyperparameter stg. In practical scenarios, to reduce computational overhead, we partition $\mathcal{D}_{\mathrm{u}}$ into multiple subsets rather than iterating over each individual data point $\boldsymbol{s}_{\mathrm{u}}$.
Algorithm 2 TRU Framework
1:Input: Initial parameters $\boldsymbol{\theta}_{\mathrm{org}}$, learning rate lr, number of iterations $T$, hyperparameter stg, and unlearning dataset $\mathcal{D}_{\mathrm{u}}$.
2:// Initialize the aggregated (rectified) task vector.
3:$\tilde{\boldsymbol{T}}_{\mathcal{D}_{\mathrm{u}}}\leftarrow\mathbf{0}$.
4:for each data point $\boldsymbol{s}_{\mathrm{u}}\in\mathcal{D}_{\mathrm{u}}$ do
5:     // 1) Compute the “targeted” task vector for $\boldsymbol{s}_{\mathrm{u}}$.
6:     $\boldsymbol{T}_{\boldsymbol{s}_{\mathrm{u}}}\;\leftarrow\;\textsc{ComputeTargetedTaskVector}\bigl{(}\boldsymbol{s}_{\mathrm{u}};\,\boldsymbol{\theta}_{\mathrm{org}}\bigr{)}$;
7:     // 2) Compute the “retain” task vector by finetuning on $\mathcal{D}_{\mathrm{u}}\setminus\{\boldsymbol{s}_{\mathrm{u}}\}$.
8:     $\boldsymbol{T}_{\mathrm{ret}}\;\leftarrow\;\textsc{ComputeRetainTaskVector}\bigl{(}\mathcal{D}_{\mathrm{u}}\setminus\{\boldsymbol{s}_{\mathrm{u}}\};\,\boldsymbol{\theta}_{\mathrm{org}}\bigr{)}$;
9:     // 3) Check for conflict: if $-\boldsymbol{T}_{\boldsymbol{s}_{\mathrm{u}}}$ negatively impacts $\boldsymbol{T}_{\mathrm{ret}}$, perform rectification.
10:     if $\langle-\boldsymbol{T}_{\boldsymbol{s}_{\mathrm{u}}},\boldsymbol{T}_{\mathrm{ret}}\rangle<0$ then
11:         $\tilde{\boldsymbol{T}}_{\boldsymbol{s}_{\mathrm{u}}}\;\leftarrow\;\boldsymbol{T}_{\boldsymbol{s}_{\mathrm{u}}}\;+\;\frac{\Bigl{[}\langle\boldsymbol{T}_{\boldsymbol{s}_{\mathrm{u}}},\,\boldsymbol{T}_{\mathrm{ret}}\rangle\Bigr{]}_{-}}{\|\boldsymbol{T}_{\mathrm{ret}}\|^{2}}\;\boldsymbol{T}_{\boldsymbol{s}_{\mathrm{u}}}\,.$
12:     else
13:         $\tilde{\boldsymbol{T}}_{\boldsymbol{s}_{\mathrm{u}}}\leftarrow\boldsymbol{T}_{\boldsymbol{s}_{\mathrm{u}}}$.
14:     end if
15:     // 4) Accumulate rectified task vectors for all points in $\mathcal{D}_{\mathrm{u}}$.
16:     $\tilde{\boldsymbol{T}}_{\mathcal{D}_{\mathrm{u}}}\;\leftarrow\;\tilde{\boldsymbol{T}}_{\mathcal{D}_{\mathrm{u}}}\;+\;\tilde{\boldsymbol{T}}_{\boldsymbol{s}_{\mathrm{u}}}$.
17:end for
18:// 5) Apply the aggregated (rectified) vector once, scaled by $\texttt{stg}/n$.
19:$\boldsymbol{\theta}_{\mathrm{TRU}}\;\leftarrow\;\boldsymbol{\theta}_{\mathrm{org}}\;-\;\frac{\texttt{stg}}{n}\;\tilde{\boldsymbol{T}}_{\mathcal{D}_{\mathrm{u}}}\,.$
20:Return $\boldsymbol{\theta}_{\mathrm{TRU}}$.

### v3

@@ Appendix A Derivations and Proofs
$\Delta$ $=\mathcal{R}(\bm{\theta}_{\mathrm{u}}^{(t+1)})\;-\;\mathcal{R}(\bm{\theta}_{\mathrm{gru}}^{(t+1)})$
$\geq\Bigl{[}\underbrace{\mathcal{R}(\bm{\theta}^{(t)})\;-\;\texttt{lr}\,\langle\bm{g}_{\mathrm{r}}^{(t)},\,\bm{g}_{\mathrm{u}}^{(t)}\rangle\;+\;\frac{\ell\,\texttt{lr}^{2}}{2}\,\|\bm{g}_{\mathrm{u}}^{(t)}\|^{2}}_{\text{Lower bound for }\mathcal{R}(\bm{\theta}_{\mathrm{u}}^{(t+1)})}\Bigr{]}\;-\;\Bigl{[}\underbrace{\mathcal{R}(\bm{\theta}^{(t)})\;-\;\texttt{lr}\,\langle\bm{g}_{\mathrm{r}}^{(t)},\,\tilde{\bm{g}}_{\mathrm{u}}^{(t)}\rangle\;+\;\frac{L\,\texttt{lr}^{2}}{2}\,\|\tilde{\bm{g}}_{\mathrm{u}}^{(t)}\|^{2}}_{\text{Upper bound for }\mathcal{R}(\bm{\theta}_{\mathrm{gru}}^{(t+1)})}\Bigr{]}.$
@@
$\Delta$ $\;\geq\;\Bigl{[}-\,\texttt{lr}\,\langle\bm{g}_{\mathrm{r}}^{(t)},\,\bm{g}_{\mathrm{u}}^{(t)}\rangle\;+\;\tfrac{\ell\,\texttt{lr}^{2}}{2}\,\|\bm{g}_{\mathrm{u}}^{(t)}\|^{2}\Bigr{]}\;-\;\Bigl{[}-\,\texttt{lr}\,\langle\bm{g}_{\mathrm{r}}^{(t)},\,\tilde{\bm{g}}_{\mathrm{u}}^{(t)}\rangle\;+\;\tfrac{L\,\texttt{lr}^{2}}{2}\,\|\tilde{\bm{g}}_{\mathrm{u}}^{(t)}\|^{2}\Bigr{]}$
$\;=\;\underbrace{-\,\texttt{lr}\,\langle\bm{g}_{\mathrm{r}}^{(t)},\,\bm{g}_{\mathrm{u}}^{(t)}\rangle\;+\;\texttt{lr}\,\langle\bm{g}_{\mathrm{r}}^{(t)},\,\tilde{\bm{g}}_{\mathrm{u}}^{(t)}\rangle}_{\text{(linear-difference term)}}\;+\;\underbrace{\frac{\ell\,\texttt{lr}^{2}}{2}\,\|\bm{g}_{\mathrm{u}}^{(t)}\|^{2}\;-\;\frac{L\,\texttt{lr}^{2}}{2}\,\|\tilde{\bm{g}}_{\mathrm{u}}^{(t)}\|^{2}}_{\text{(quadratic-difference term)}}.$
@@
    Rectification Nonnegativity. Since $\tilde{\bm{g}}_{\mathrm{u}}^{(t)}$ is formed from $\bm{g}_{\mathrm{u}}^{(t)}$ by removing negatively aligned components with respect to $\bm{g}_{\mathrm{r}}^{(t)}$, we have $\langle\bm{g}_{\mathrm{r}}^{(t)},\,\tilde{\bm{g}}_{\mathrm{u}}^{(t)}\rangle\geq\langle\bm{g}_{\mathrm{r}}^{(t)},\,\bm{g}_{\mathrm{u}}^{(t)}\rangle$, and thus $-\,\langle\bm{g}_{\mathrm{r}}^{(t)},\,\bm{g}_{\mathrm{u}}^{(t)}\rangle\;+\;\langle\bm{g}_{\mathrm{r}}^{(t)},\,\tilde{\bm{g}}_{\mathrm{u}}^{(t)}\rangle\geq 0$. Multiplication by lr preserves non-negativity, ensuring that the expression inside the first bracket remains non-negative.
@@
    Curvature condition.  By construction $\|\tilde{\bm{g}}_{\mathrm{u}}^{(t)}\|=\|\bm{g}_{\mathrm{u}}^{(t)}\|\sin\phi$, where $\phi$ is the angle between $\bm{g}_{\mathrm{u}}^{(t)}$ and $\bm{g}_{\mathrm{r}}^{(t)}$. Condition a) of the theorem states $\ell\geq L\!\left(1-\cos^{2}\phi\right)=L\sin^{2}\phi$. Therefore
[…]
    implying the quadratic‐difference term is non–negative for every $0<\texttt{lr}\leq 2/L$ (condition b)).
Since both terms are non–negative, we have $\Delta\geq 0$, i.e., $\mathcal{R}(\bm{\theta}^{(t+1)}_{\mathrm{gru}})\leq\mathcal{R}(\bm{\theta}^{(t+1)}_{\mathrm{u}})$. ∎
@@ Appendix B Detailed Results
This section provides comprehensive results that echo our main experiments discussed in Section 5.1. It encompasses TOFU, WMDP, and MUSE benchmarks, further incorporating additional baseline methods like SimNPO, and other metrics, such as PrivLeak for MUSE. These results are summarized in Tables 2-4.
@@
Appendix C Comparison with RMU on WMDP
For completeness, we compare our approach with RMU, the method proposed alongside WMDP. Due to sensitivity issues noted in the official implementation, we set the hyperparameter $\alpha$ to 100 (instead of the default 1200) to ensure stable optimization. Table 5 reports the results for both RMU and its combination with GRU.
Table 5: Comparison of RMU and RMU with GRU on the WMDP benchmark (Bio and Cyber: accuracy $\downarrow$; MMLU: accuracy $\uparrow$).
Method Bio $\downarrow$ Cyber $\downarrow$ MMLU $\uparrow$
RMU 0.26 0.31 0.41
w/ GRU 0.26 0.28 0.44
As shown, GRU consistently improves both unlearning and retention over the RMU baseline.
Appendix D Hyper-parameter Analyses
In addition to our main results, we further discuss about our hyper-parameter configurations as well as conduct additional analyses on hyper-parameter sensitivity.
D.1 Hyper-parameter Configurations
We employ grid search on validation data to select proper hyper-parameters for GRU and TRU. For GRU, the candidate values for $\gamma$ include $\{0.01,0.05,0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,0.99\}$, and that for $\tau$ are $\{0.001,0.005,0.01,0.1,1.0,10,100\}$. For TRU, we select stg from the space $\{0.50,0.60,0.65,0.70,0.75,0.80,0.85\}$ while fix $\texttt{lr}=1e^{-4}$. We further summarize their detailed configurations across different setups as follows.
@@ D.1 Hyper-parameter Configurations
D.2 Sensitivity Analyses
Method Metric $\tau$
0.001 0.01 0.1 1.0 2.0 3.0 10 100 N/A
Method Metric $\gamma$
0.01 0.02 0.05 0.1 0.2 0.3 0.4 0.5 0.6 0.8 0.9 0.99 N/A
Appendix E Ablation Studies and Other Analyses
@@
Ablation Studies. Previous works, such as (Zhang et al., 2024), also use gradient clipping (GC) to improve the overall efficacy of unlearning, raising us to ask if our rectification mechanism plays a key role to mitigate the trade-off between removal and retention. In Table 8, we conduct ablation studies on TOFU using Llama-2-7B as the backbone, focusing on the 5% unlearning setup. We compare three scenarios: the original unlearning method without GRU (w/o GRU), the original method enhanced with GC (w/ GC), and the unlearning method that incorporates GRU (w/ GRU). As evident from the results, GRU demonstrates superior scores in terms of both FQ and MU, showing its efficacy in mitigating the trade-off between removal and retention.
Method Component FQ $\uparrow$ MU $\uparrow$
GA w/o GRU -16.9281 0.0000
w/ GC -20.7646 0.0000
w/ GRU -3.2299 0.5149
NPO w/o GRU -10.9105 0.4919
w/ GC -10.9105 0.4970
w/ GRU -10.4275 0.5373
@@
Visualization of Rectification. We further examine the angles between $\bm{g}_{\mathrm{u}}^{(t)}$ and $\bm{g}_{\mathrm{r}}^{(t)}$, along with those results after being rectified via GRU. We monitor the dynamics of these angles throughout the unlearning processes for various baseline methods, as well as the changes after applying GRU for rectification. As a case study, Figure LABEL:fig:cos_theta_comparison shows these results on TOFU 5% unlearning with Llama-2-7B as the backbone. Without the use of GRU, the cosine similarity between $\bm{g}_{\mathrm{u}}^{(t)}$ and $\bm{g}_{\mathrm{r}}^{(t)}$ keeps negative throughout unlearning, suggesting potential adverse effects on the overall performance of the model. In comparison, within the unlearning dynamics facilitated by GRU, it is observed that although the angles initially continue to be negative (dotted lines), our rigorous method of gradient rectification will adjust the resulting cosine similarity to exactly 0. This adjustment ensures that the gradient direction associated with unlearning is completely orthogonal to that of retention, thereby effectively maintaining the overall model performance.
Appendix F Aligning Unlearning with Retention: A Practically Meaningful Evaluation
Recalling the dual goals of LLM unlearning, a meaningful evaluation requires not only improvements in the forget set, but also that the retention utility of the model remains well aligned with its original capabilities. A substantial decline in utility performance would render the resulting model ineffective, making the process of LLM unlearning itself meaningless. This concern, often overlooked in recent unlearning studies, can be addressed by Unlearning with Control (UWC) (Wang et al., 2024a), as discussed in Section 2.2. UWC provides a post-unlearning calibration framework that restores retention performance by interpolating the model parameters before and after unlearning via a tunable parameter $\alpha$. Leveraging UWC calibration, we systematically evaluate our approach under explicit retention thresholds (e.g., 85%, 90%, and 95%) to investigate whether strong unlearning performance can be achieved without sacrificing essential retention utility, thus aligning unlearning objectives with practical deployment needs. Given the similarity of our findings across multiple benchmarks, we focus here on representative results obtained from the challenging Phi setup on TOFU. Specifically, we present GA and NPO as baseline methods to illustrate the flexibility of UWC calibration and highlight the effectiveness of our GRU approach in attaining superior unlearning results while rigorously maintaining retention utility. The results in Tables 9 and 10 clearly demonstrate that, under each retention constraint, incorporating GRU consistently yields substantial improvements in forget quality (FQ) compared to the calibrated baselines, while perfectly maintaining the prescribed model utility (MU). This pattern holds for both GA and NPO methods, across all retention targets and unlearning setups. These findings validate the practical value of our approach: by leveraging UWC calibration in combination with GRU, practitioners can achieve strong, controllable unlearning effects without sacrificing the essential utility of large language models, thereby ensuring that unlearning objectives remain aligned with real-world deployment requirements.
Table 9: GA on the TOFU Phi-1.5 setup with UWC calibration. We report FQ (forget quality, $\uparrow$) and MU (model utility, $\uparrow$) for 5% and 10% unlearning under three retention targets (85%, 90%, 95%). Each retention target is shown in two adjacent rows: the first row gives the calibrated GA result, and the second (w/ GRU) shows the result after incorporating GRU. The better score within each GA–GRU pair is in bold.
Method 5% Unlearning 10% Unlearning
[…]
Original -28.8 0.52 -40.5 0.52
GA (85%) -22.0 0.44 -35.3 0.44
w/ GRU (85%) -8.6 0.44 -20.8 0.44
GA (90%) -28.1 0.47 -36.8 0.47
w/ GRU (90%) -15.2 0.47 -28.8 0.47
GA (95%) -28.1 0.49 -39.8 0.49
w/ GRU (95%) -18.8 0.49 -33.9 0.49
Table 10: NPO on the TOFU Phi-1.5 setup with UWC calibration. Layout and notation follow Table 9.
Method 5% Unlearning 10% Unlearning
[…]
Original -28.8 0.52 -40.5 0.52
NPO (85%) -15.7 0.44 -31.9 0.44
w/ GRU (85%) -9.5 0.44 -15.2 0.44
NPO (90%) -20.1 0.47 -35.3 0.47
w/ GRU (90%) -12.9 0.47 -18.2 0.47
NPO (95%) -25.0 0.49 -38.2 0.49
w/ GRU (95%) -14.0 0.49 -20.8 0.49
Appendix G Comparison with Gradient Direction Rectification (GDR)
Closely related is Gradient Direction Rectification (GDR) (Lin et al., 2024), as it similarly employs gradient projection to resolve conflicts between forgetting and retention objectives. However, GDR relies on caching gradients across epochs, resulting in substantial memory overhead that limits its scalability for large language models, where parameter sizes are massive and training typically involves only a few epochs. In contrast, GRU dynamically estimates retention gradients using an exponential moving average, greatly reducing memory cost and enabling practical unlearning at scale. Furthermore, while GDR merges retention gradients directly into parameter updates, potentially increasing the risk of overfitting to the retention set, our GRU approach leverages retention gradients solely as constraints for rectification. Finally, TRU extension addresses the challenge of biased retain data, a scenario unique to LLM unlearning and not considered in GDR.